# Learning Informative Latent Representation for Quantum State Tomography

### v1

@@ Abstract
1 Introduction
The determination of the quantum state of a system, known as quantum state tomography (QST), is the gold standard for verification and benchmarking of quantum devices and is also essential for engineering systems in developing quantum technology [14, 39]. In the physical world, the full description of a quantum state can be represented as a density matrix $\rho$, i.e., a positive-definite Hermitian matrix with unit trace. To determine $\rho$, one may first perform measurements on a collection of identically prepared copies of a quantum system, gathering statistical outcomes to a set of frequencies $\mathbf{f}$, as shown in Figure 0(a). Then, the reconstruction of quantum states is realized by mapping the measured frequencies $\mathbf{f}$ into a full description of a quantum state (density matrix $\rho$) or a partial description of a quantum state (quantum properties $\bm{\mu}$) using approximation algorithms, as illustrated in Figure 0(b).
To uniquely identify a quantum state, the measurements must be informatively complete to provide all the information about $\rho$ [21]. The exponential scaling of parameters in $\rho$ requires an exponentially increasing number of measurements, each of which requires a sufficient number of identical copies [14], posing a great challenge in practical applications. For example, in solid-state systems, the process associated with measuring one copy of a quantum state can be time-consuming, and implementing a sufficient number of measurement operators requires complex and costly experimental setups. Given a coherence time (beyond which quantum states might change), the total copies of identical states for measurements may be constrained. In this scenario, the measurement data may be a complete but inaccurate frequency vector (a few copies assigned to each measurement operator ) or an accurate but incomplete frequency vector [24] (sufficient copies assigned to partial measurement operators that are experimentally easy to generate [6]). The two factors are collectively referred to as imperfect scenarios (i.e., ill-posed problems) in this paper, which hinder the precise reconstruction of a quantum state.
@@ 1 Introduction
When faced with imperfect statistics obtained in practical scenarios, traditional methods either treat those imperfections as noise in accurately measured frequencies, without employing a specific methodology [21] or require manually-designed regularizers [40] to handle the ill-posed problem. Recently, neural network (NN) approaches have demonstrated their generality and robustness in learning intricate features from quantum systems [14]. However, these methods have not yet incorporated imperfect factors into their network design and thus exhibit sub-optimal performance when confronted with imperfect measurement data. Practical QST suffers from noise addition and information loss in the measured frequencies, reducing the representation of measured data for accurate quantum state reconstruction. With the emergence of new architectures of NNs designed for practical problems [41, 9], a natural question arises: is it possible to develop an enhanced architecture that corporates imperfect measurement data to learn an informative representation from imperfect measurement data?
@@
Recognizing the representation reduction posed by the ill-posed problem, we turn to a powerful neural architecture to extract a latent representation that might compensate for imperfect data. Specifically, we propose a transformer-based autoencoder to address the challenge of imperfect measurement data in practical QST applications. The encoder serves as an extractor to capture the informative latent representation (ILR) of quantum states from measured data. Through the state decoder, the latent representation then provides a useful feature that can be further leveraged to estimate a quantum state, including its density matrix and significant properties. However, simply training the autoencoder model with raw data might not entirely capitalize on the potential of intermediate layers to abstract meaningful latent representations from the imperfect data. Therefore, we further design a pre-training strategy that guides the encoder to learn a highly informative latent representation by augmenting a frequency decoder. The decoder retrieves the true probabilities of a complete set of measurements that provide more comprehensive information about quantum states. To ensure that the encoder’s latent representation is sufficiently informative while maintaining model expressiveness predominantly in the encoder, the pre-training autoencoder is designed with an asymmetrical architecture, where the encoder is larger in scale compared to the decoder. The entire design facilitates the acquisition of an informative latent representation from imperfect measurement data, which is highly beneficial for the reconstruction of quantum states under imperfect conditions. The key contributions of our work are summarized as follows:
@@
    Numerical simulations demonstrate that ILR has significant potential in reconstructing density matrices and predicting quantum properties. Experimental implementations on IBM quantum machines demonstrate the potential of ILR in compensating for noise in real quantum devices when directly using the model trained with simulated data.
2 Related work
@@
Traditional QST. Various approaches have been proposed for estimating quantum states based on measured statistics [35]. One commonly used method is the least-squares inversion that solves the inverse of linear equations which relates the measured quantities to the density matrix elements [32]. Maximum likelihood estimation (MLE) is another popular method that searches for the state estimate by maximizing the probability of the observed data, but it involves a large number of nonlinear equations [21]. Linear regression estimation (LRE) transforms the problem into a linear model that can be efficiently solved but requires additional physical projection techniques to avoid non-physical quantum states [35, 36]. Bayesian tomography constructs a state using an integral averaging over all possible quantum states with proper weights [20]. While compressed sensing can estimate low-rank states (with a few eigenvectors) under an incomplete measurement setting [12], determining specific measurement operators restricts its applicability in real-world scenarios.
@@
NN-based QST. QST involves extracting useful information from measured statistics [21] and has been investigated using different neural networks. For instance, fully connected networks have been applied to approximate a function that maps measured statistics into quantum parameters [42] and has been applied to denoise the state-preparation-and-measurement error [34]. A convolutional neural network has been introduced to reconstruct 2-qubit quantum states from an image of measured outcomes [27, 28]. Attention models have also been explored to capture long correlations among quantum data [5, 43, 44]. While NNs have been applied to handle incomplete measurements [7], the method relies on a separate statistical technique to infer missing values from existing data and the decoupling of various functionalities into different models. The capability of NNs has also been demonstrated by their application in predicting statistics of new measurements [45]. Here, we design an enhanced NN architecture to learn an informative latent representation from imperfect measurement data, which benefits the reconstruction of quantum systems at different levels. Compared with classical shadow tomography which is specifically designed for predicting quantum properties [19], our approach offers a unified framework that can be utilized to reconstruct density matrices and predict quantum properties. Autoencoder. Autoencoding is a neural network technique to learn latent representations through an encoder that accepts raw input data and a decoder that reconstructs the input from the latent representation [17]. It is a fundamental and powerful tool in deep learning with a wide range of applications, such as image denoising [22, 18] and reconstruction [10], natural language processing [4], and others. The transformer architecture’s self-attention mechanism [41] and pre-training technique, suitable for large dataset tasks, have popularized autoencoders in natural language processing (NLP) [9, 38, 37] and computer vision [11, 26, 15], showing superiority to convolutional neural networks (CNNs) [23, 3]. Our method employs the transformer-based autoencoder with a pre-training technique to reconstruct quantum states using the intermediate latent extracted from imperfect data.
3 Method
@@
In this section, we first introduce the basic concepts of QST and then present the ill-posed challenge for QST with imperfect measurement data. Finally, a transformer-based autoencoder architecture with a pre-training strategy is proposed to learn the informative latent representation for QST.
3.1 Preliminaries about QST
@@
QST is a process of extracting useful information about a quantum state based on a set of measurements. For a $d$-dimensinal quantum system, the full mathematical representation of a quantum state can be described as a density matrix $\rho\in\mathbb{C}^{d\times d}$. Measurement operators are usually positive-operator-valued measurements (POVMs) and can be represented as a set of positive semi-definite Hermitian matrices $\mathbb{O}_{i=1}^{M}\in\mathbb{C}^{d\times d}$ that sum to identity ($\sum_{i=1}^{M}\mathbb{O}_{i}=\mathbb{I}^{d\times d}$[31]. According to Born’s rule, when measuring the quantum state $\rho$ with the measurement operator $\mathbb{O}_{i}$, the true probability $p_{i}\in[0,1]$ of obtaining the outcome $i$ is calculated as
@@
In practice, a finite number of identical copies of a quantum state are used for measurements. In Figure 0(a), the statistical frequency for the $i$-outcome $f_{i}\in[0,1]$ is collected as
@@
where $N\in\mathbb{R}$ denotes the total number of copies of identical states and $n_{i}\in\mathbb{R}$ denotes the occurrences for the outcome $i$. As the true probability $p_{i}$ is directly inaccessible from experimental results, a natural solution is to leverage the measured frequency vector $\mathbf{f}=[f_{1},f_{2},...,f_{M}]^{T}\in\mathbb{R}^{M\times 1}$ which is a statistical approximation to the true probability vector $\mathbf{p}=[p_{1},p_{2},...,p_{M}]^{T}\in\mathbb{R}^{M\times 1}$, to infer underlying information about the quantum state $\rho$. As illustrated in Figure 0(b), reconstructing quantum states can be solved by approximating a function that maps the frequency vector $\mathbf{f}$ into desired representations of a quantum state, e.g., a density matrix $\rho$. In addition, it is also useful to deduce significant properties such as purity, entanglement, and entropy [19], which can be denoted as a real vector $\bm{\mu}=[\mu_{1},\mu_{2},...,\mu_{k}]^{T}\in\mathbb{R}^{k\times 1}$, where $k$ denotes the number of properties.
3.2 The ill-posed QST problem
@@
The imperfect factors discussed above correspond to an ill-posed problem, which is a common issue in machine learning. In the context of quantum estimation tasks, ill-posedness results in an inaccurate or incomplete observation of quantum states, leading to errors in subsequent QST tasks. To provide a clear explanation, we conduct an initial experiment on 2-qubit pure states using the simplest fully connected network (FCN) [29] and a traditional linear regression estimation LRE [35]. The log of infidelity is measured as $\bar{F}=\log(1-F(\rho,\hat{\rho}))\in\mathbb{R}$, where $F(\rho,\hat{\rho})=|\textup{Tr}(\sqrt{\sqrt{\rho}\hat{\rho}\sqrt{\rho}})|$ denotes the fidelity between the reconstructed state $\hat{\rho}$ and the original state $\rho$ [31]. The numerical results in Figure 2 show a strong correlation between the infidelity and the copies of the identical states or the completeness of the measurement. Specifically, we observe a sharp decrease in infidelity as the number of copies increases, while masking $8$ operators among the original complete measurements results in a large increase in infidelity. The underlying reasons might be that the noise addition and information loss in the imperfect measurement data result in a diminished representation of quantum states. The observed ill-posed challenge promotes the necessity of acquiring an informative latent representation that can potentially compensate for the imperfections in practical QST settings.
3.3 Learning an informative latent representation for QST with imperfect measurement data
@@
Given the representation reduction posed by ill-posed QST, it is imperative to design a robust neural architecture capable of learning a latent representation that might compensate for imperfect measurement data. Taking inspiration from the success of transformer-based autoencoders in learning latent representations from imperfect data in NLP and CV tasks, we propose an autoencoder to learn a latent representation from imperfect measurement data. The latent representation provides an informative feature that can be further leveraged to estimate quantum states by a decoder, including its density matrix and significant properties. To further enhance the informative latent representation, a pre-training process is implemented to retrieve high-quality data from imperfect measurement data by augmenting a frequency decoder. The schematic of our approach is illustrated in Figure 3.
@@
Pre-training process using a masked and asymmetrical autoencoder. The transformer-based encoder utilized in the QST process is specifically designed to be expressive and proficient in extracting latent features from imperfect data. However, simply training the autoencoder model with raw data may not fully utilize the potential of the intermediate layers to abstract an informative latent representation from the imperfect measurement data. A high-dimensional ILR has the capability to capture a broader range of information pertaining to quantum states, resulting in enhanced accuracy during the reconstruction process (see Figure 2). To obtain an improved ILR for QST with imperfect measurement data, we employ a pre-training strategy to guide the encoder to retrieve perfect measured data, i.e., true probabilities of a complete set of measurements. We reuse the transformer-based encoder from the QST process and combine it with a lightweight frequency decoder to approximate the true probabilities from imperfect frequencies.
@@
Training strategies. The pre-training process is first implemented to obtain a highly representative representation of quantum states. Then, complete training of the QST process is achieved by keeping the parameters of the transformer-encoder fixed and fine-tuning the parameters of the state decoder. This structure enables the preservation of the information that is learned from the pre-trained model, and the latent representation enables the state decoder to reconstruct quantum states with improved performance. To simulate different incomplete measurement scenarios, we introduce two ways for pre-training: i) Separate, fixed mask values are utilized for all samples, and ii) Unified, where mask values are randomly selected for each training batch. Similarly, the QST process has the same two training options. Please refer to Appendix B for their algorithmic descriptions. Now, there are four different strategies for training the whole model, including S2S (separate pre-training with separate tomography), S2U (separate pre-training with unified tomography), U2S (unified pre-training with separate tomography), and U2U (unified pre-training with unified tomography). We conduct a comprehensive comparison of 2-qubit pure states and determine S2U as a good strategy that balances accuracy and efficiency. Please refer to Appendix B for detailed information.
Remark 1
@@
One operator $\mathbb{O}_{i}\in\mathbb{C}^{d\times d}$ can be rewritten as a real column vector $O_{j}\in\mathbb{R}^{2d^{2}\times 1}$ by splitting its real and imaginary parts. A set of operators $\{\mathbb{O}_{1},\mathbb{O}_{2},...,\mathbb{O}_{M}\}$ can be concatenated into an operator matrix $\mathbf{O}=[O_{1},O_{2},...,O_{M}]\in\mathbb{R}^{M\times 2d^{2}}$.
Remark 2
@@
Generate a property vector $\bm{\mu}$: Let the number of properties be $K$, and then multiple properties form a real vector $\bm{\mu}\in\mathbb{R}^{K\times 1}$, with the ground-truth for each element calculated from the density matrix $\rho$ (Detailed information provided in Appendix A).
Remark 3

### v2

@@ Abstract
Index Terms:
quantum state tomography, imperfect measurement data, fidelity, latent representation, pre-training.
I Introduction
The determination of the quantum state of a system, known as quantum state tomography (QST), is the gold standard for verification and benchmarking of quantum devices and is also essential for engineering systems in developing quantum technology [1, 2]. In the physical world, the full description of a quantum state can be represented as a density matrix $\rho$, i.e., a positive-definite Hermitian matrix with unit trace. To determine $\rho$, one may first perform measurements on a collection of identically prepared copies of a quantum system, gathering statistical outcomes to a set of frequencies $\mathbf{f}$, as shown in Fig. 1(a). Then, the reconstruction of quantum states is realized by mapping the measured frequencies $\mathbf{f}$ into a full description of a quantum state (density matrix $\rho$) or a partial description of a quantum state (quantum properties $\bm{\mu}$) using approximation algorithms, as illustrated in Fig. 1(b).
Generally, to uniquely identify a quantum state, the measurements must be informatively complete to provide all the information about $\rho$ [3], except for some special cases when density matrices of quantum states are low-rank [4] or take the form of matrix product operators [5]. The exponential scaling of parameters in $\rho$ requires an exponentially increasing number of measurements, each of which requires a sufficient number of identical copies [1], posing a great challenge in practical applications. For example, in solid-state systems, the process associated with measuring one copy of a quantum state can be time-consuming, and implementing a sufficient number of measurement operators requires complex and costly experimental setups. Given a coherence time (beyond which quantum states might change), the total copies of identical states for measurements may be constrained. In this scenario, the measurement data may be a complete but inaccurate frequency vector (a few copies assigned to each measurement operator ) or an accurate but incomplete frequency vector [6] (sufficient copies assigned to partial measurement operators that are experimentally easy to generate [7]). The two factors are collectively referred to as imperfect scenarios (i.e., ill-posed problems) in this paper, which hinder the precise reconstruction of a quantum state.
@@ I Introduction
When faced with imperfect statistics obtained in practical scenarios, traditional methods either treat those imperfections as noise in accurately measured frequencies, without employing a specific methodology [3] or require manually-designed regularizers [8] to handle the ill-posed problem. Recently, neural network (NN) approaches have demonstrated their generality and robustness in learning intricate features from quantum systems [1], including restricted Boltzmann machine (RBM) [9] and recurrent neural network (RNN) [10, 11] and generative adversarial network (GAN) [12] for QST tasks. However, these methods have not yet incorporated imperfect factors into their network design and thus exhibit sub-optimal performance when confronted with imperfect measurement data. Practical QST suffers from noise addition and information loss in the measured frequencies, reducing the representation of measured data for accurate quantum state reconstruction. Although the capability of neural networks in dealing with incomplete measurement has been investigated in [13], their method relies on a separate statistical technique to infer missing values from the rest of the data and involves the decoupling of various functionalities gathered in one single model into different models. With the emergence of new architectures of NNs designed for practical problems [14, 15], a natural question arises: is it possible to develop an enhanced architecture that corporates imperfect measurement data to learn an informative representation from imperfect measurement data?
@@
Recognizing the representation reduction posed by the ill-posed problem, we turn to a powerful neural architecture to extract a latent representation that might compensate for imperfect data. Specifically, we propose a transformer-based autoencoder to address the challenge of imperfect measurement data in practical QST applications. The encoder serves as an extractor to capture the informative latent representation (ILR) of quantum states from measured data. Through the state decoder, the latent representation then provides a useful feature that can be further leveraged to estimate a quantum state, including its density matrix and significant properties. In typical ML tasks, supervised training, by its nature, is a guided process. When it relies solely on imperfect data as the ground truth, it restricts the intermediate layers from fully realizing their potential, thereby limiting the richness of the latent representations. Hence, simply training the autoencoder model with raw measured data might not entirely capitalize on the potential of intermediate layers to abstract meaningful latent representations from the imperfect data. To overcome the limitation of imperfect measured data, we propose augmenting the encoder with a frequency decoder during pre-training, which helps to extract a highly informative latent representation.
To ensure that the encoder’s latent representation is sufficiently informative while maintaining model expressiveness predominantly in the encoder, the pre-training autoencoder is designed with an asymmetrical architecture, where the encoder is larger in scale compared to the decoder. The entire design facilitates the acquisition of an informative latent representation from imperfect measurement data, which is highly beneficial for the reconstruction of quantum states under imperfect conditions. The key contributions of our work are summarized as follows:
@@
    Numerical simulations demonstrate that ILR has significant potential in reconstructing density matrices over other NN-based methods. Experimental implementations on IBM quantum machines demonstrate the potential of ILR in compensating for noise in real quantum devices when directly using the model trained with simulated data.
II Related work
In this section, we first review traditional QST methods and then introduce the NN-based QST method. Finally, we review transformer-based autoencoders.
II-A Traditional QST
@@
Various approaches have been proposed for estimating quantum states based on measured statistics [16]. One commonly used method is the least-squares inversion that solves the inverse of linear equations which relates the measured quantities to the density matrix elements [17]. Maximum likelihood estimation (MLE) is another popular method that searches for the state estimate by maximizing the probability of observed data, but it involves a large number of nonlinear equations [3]. Linear regression estimation (LRE) transforms the problem into a linear model that can be efficiently solved but requires additional physical projection techniques to avoid non-physical quantum states [16, 18]. Bayesian tomography constructs a state using an integral averaging over all possible quantum states with proper weights [19]. While compressed sensing can estimate low-rank states under an incomplete measurement setting [20], determining specific measurement operators restricts its applicability in real-world scenarios. When focusing on estimating partial properties, classical shadow estimation [21] and randomized benchmarking [22] are useful tools owing to their polynomial scale with system size.
II-B NN-based QST
@@
QST involves extracting useful information from measured statistics [3] and has been investigated using different neural networks. For instance, fully connected networks have been applied to approximate a function that maps measured statistics into quantum parameters [23] and has been applied to denoise the state-preparation-and-measurement error [24]. A convolutional neural network has been introduced to reconstruct 2-qubit quantum states from an image of measured outcomes [25, 26]. Attention models have also been explored to capture long correlations among quantum data [11, 27, 28]. While NNs have been applied to handle incomplete measurements [13], the method relies on a separate statistical technique to infer missing values from existing data and the decoupling of various functionalities into different models. The capability of NNs has also been demonstrated by their application in reconstructing density matrices [12] and predicting statistics of new measurements [29]. Compared to their methods utilizing a generative network, we design an enhanced NN architecture to learn an informative latent representation from imperfect measurement data, which benefits the reconstruction of quantum systems at different levels. Compared with classical shadow tomography which is specifically designed for predicting quantum properties [21], our approach offers a unified framework that can be utilized to reconstruct density matrices and predict quantum properties. Compared to other works that utilize NNs for different tasks [30, 9, 31], our approach aims to learn an informative representation, which is shared between different tasks.
II-C Transformer-based Autoencoder
Autoencoding is a neural network technique to learn latent representations through an encoder that accepts raw input data and a decoder that reconstructs the input from the latent representation [32]. It is a fundamental and powerful tool in deep learning with a wide range of applications, such as image denoising [33, 34] and reconstruction [35], natural language processing [36], and others. The transformer architecture’s self-attention mechanism [14] and pre-training technique, suitable for large dataset tasks, have popularized autoencoders in natural language processing (NLP) [15, 37, 38] and computer vision [39, 40, 41], showing superiority to convolutional neural networks (CNNs) [42, 43]. Inspired by the potential of attention layers to capture long-range correlation among their constituent qubits [11, 28], we design a transformer-based autoencoder with a pre-training technique to reconstruct quantum states using the intermediate latent extracted from imperfect data.
III Method
@@
In this section, we first introduce the basic concepts of QST and then present the ill-posed challenge for QST with imperfect data. Finally, a transformer-based autoencoder architecture with a pre-training strategy is proposed to learn the informative latent representation for QST.
III-A Preliminaries about QST
@@
QST is a process of extracting useful information about a quantum state based on a set of measurements. For a $n$-qubit quantum system with dimension $d=2^{n}$, the full mathematical representation of a quantum state can be described as a density matrix $\rho\in\mathbb{C}^{d\times d}$. Measurement operators are usually positive-operator-valued measurements (POVMs) and can be represented as a set of positive semi-definite Hermitian matrices $\mathbb{O}_{i=1}^{M}\in\mathbb{C}^{d\times d}$ that sum to identity ($\sum_{i=1}^{M}\mathbb{O}_{i}=\mathbb{I}^{d\times d}$[44]. According to Born’s rule, when measuring the quantum state $\rho$ with the measurement operator $\mathbb{O}_{i}$, the true probability $p_{i}\in[0,1]$ of obtaining the outcome $i$ is calculated as
@@
In practice, a finite number of identical copies of a quantum state are used for measurements. In Fig. 1 (a), the statistical frequency for the $i$-outcome $f_{i}\in[0,1]$ is collected as
@@
where $N\in\mathbb{R}$ denotes the total number of copies of identical states and $n_{i}\in\mathbb{R}$ denotes the occurrences for the outcome $i$. As the true probability $p_{i}$ is directly inaccessible from experimental results, a natural solution is to leverage the measured frequency vector $\mathbf{f}=[f_{1},f_{2},...,f_{M}]^{T}\in\mathbb{R}^{M\times 1}$ which is a statistical approximation to the true probability vector $\mathbf{p}=[p_{1},p_{2},...,p_{M}]^{T}\in\mathbb{R}^{M\times 1}$, to infer underlying information about the quantum state $\rho$. As illustrated in Fig. 1 (b), reconstructing quantum states can be solved by approximating a function that maps the frequency vector $\mathbf{f}$ into desired representations of a quantum state, e.g., a density matrix $\rho$. In addition, it is also useful to deduce significant properties such as purity, entanglement, and entropy [21], which can be denoted as a real vector $\bm{\mu}=[\mu_{1},\mu_{2},...,\mu_{k}]^{T}\in\mathbb{R}^{k\times 1}$, where $k$ denotes the number of properties.
III-B The ill-posed QST problem
@@
The imperfect factors discussed above correspond to an ill-posed problem, which is a common issue in machine learning. In the context of quantum estimation tasks, ill-posedness results in an inaccurate or incomplete observation of quantum states, leading to errors in subsequent QST tasks. To provide a clear explanation, we conduct an initial experiment on 2-qubit pure states using the simplest fully connected network (FCN) [46] and a traditional linear regression estimation LRE [16]. The log of infidelity is measured as $\bar{F}=\log_{10}(1-F(\rho,\hat{\rho}))\in\mathbb{R}$, where $F(\rho,\hat{\rho})=|\textup{Tr}(\sqrt{\sqrt{\rho}\hat{\rho}\sqrt{\rho}})|$ denotes the fidelity between the reconstructed state $\hat{\rho}$ and the original state $\rho$ [44]. The numerical results in Fig. 2 show a strong correlation between the infidelity and the copies of the identical states or the completeness of the measurement. Specifically, we observe a sharp decrease in infidelity as the number of copies increases, while masking $8$ operators among the original complete measurements (a total of 36 measurements) results in a large increase in infidelity. The underlying reasons might be that the noise addition and information loss in the imperfect measurement data result in a diminished representation of quantum states. The observed ill-posed challenge promotes the necessity of acquiring an informative latent representation that can potentially compensate for the imperfections in practical QST settings.
III-C Learning an informative latent representation for QST with imperfect measurement data
@@
Given the representation reduction posed by ill-posed QST, it is imperative to design a robust neural architecture capable of learning a latent representation that might compensate for imperfect measurement data. Taking inspiration from the success of transformer-based autoencoders in learning latent representations from imperfect data in NLP and CV tasks, we propose an autoencoder to learn a latent representation from imperfect measurement data. The latent representation provides an informative feature that can be further leveraged to estimate quantum states by a decoder, including its density matrix and significant properties. To further enhance the informative latent representation, a pre-training process is implemented to retrieve high-quality data from imperfect measurement data by augmenting a frequency decoder. The schematic of our approach is illustrated in Fig. 3.
@@
Pre-training process using a masked and asymmetrical autoencoder. The transformer-based encoder utilized in the QST process is specifically designed to be expressive and proficient in extracting latent features from imperfect data. However, simply training the autoencoder model with raw data may not fully utilize the potential of the intermediate layers to abstract an informative latent representation from the imperfect measurement data. A high-dimensional ILR has the capability to capture a broader range of information pertaining to quantum states, resulting in enhanced accuracy during the reconstruction process (see Fig. 2). To obtain an improved ILR for QST with imperfect measurement data, we employ a pre-training strategy to guide the encoder to retrieve perfect measured data, i.e., true probabilities of a complete set of measurements. We reuse the transformer-based encoder from the QST process and combine it with a lightweight frequency decoder to approximate the true probabilities from imperfect frequencies.
@@
Remark 1.
@@
One operator $\mathbb{O}_{i}\in\mathbb{C}^{d\times d}$ can be rewritten as a real column vector $O_{j}\in\mathbb{R}^{2d^{2}\times 1}$ by splitting its real and imaginary parts. A set of operators $\{\mathbb{O}_{1},\mathbb{O}_{2},...,\mathbb{O}_{M}\}$ can be concatenated into a matrix $\mathbf{O}=[O_{1},O_{2},...,O_{M}]\in\mathbb{R}^{M\times 2d^{2}}$.
Remark 2.
@@
Generate a property vector $\bm{\mu}$: Let the number of properties be $K$, and then multiple properties form a real vector $\bm{\mu}\in\mathbb{R}^{K\times 1}$, with the ground-truth for each element calculated from the density matrix $\rho$ (Detailed information provided in Appendix A-A).
Remark 3.